### v1

@@ 6.1 Individual polychoric correlation coefficient
For misspecification fraction $\varepsilon\in\{0,0.1,0.2\}$, we sample $N=1,000$ ordinal responses from this data generating process and estimate the true parameter $\bm{\theta}_{*}$ with the MLE (Olsson,, 1979), our proposed estimator with tuning constant set to $c=1.6$, and, for comparison, the Pearson sample correlation calculated on the integer-valued responses. We repeat this procedure for 1,000 simulated datasets. As performance measures, we calculate the average bias, standard deviation across repetitions, coverage, and length of confidence intervals at significance level $\alpha=0.05$. The coverage is defined as proportion of $(1-\alpha)$-th confidence intervals $[\widehat{\rho}_{N}\mp q_{1-\alpha/2}\cdot\text{SE}(\widehat{\rho}_{N})]$ that contain the true $\rho_{*}$, where $q_{1-\alpha/2}$ is the $(1-\alpha/2)$-th quantile of the standard normal distribution and $\text{SE}(\widehat{\rho}_{N})$ is the standard error of $\widehat{\rho}_{N}$, which is constructed using the limit theory developed in Theorem 1 in Appendix A. The length of a confidence interval is given by $2\cdot q_{1-\alpha/2}\cdot\text{SE}(\widehat{\rho}_{N})$.
Figure 4: Boxplot visualization of the bias of three estimators of the polychoric correlation coefficient, $\widehat{\rho}_{N}-\rho_{*}$, for various degrees of misspecification of the polychoric model across 1,000 replications. The estimators are the proposed robust estimator with $c=1.6$ (left), the MLE of Olsson, (1979) (center), and the Pearson sample correlation coefficient (right).
@@
Table 1: Performance measures of of the three estimators, Pearson sample correlation, the MLE, and the proposed robust estimator with $c=1.6$, for various misspecification fractions across 1,000 simulated datasets. The true polychoric correlation coefficient is $\rho_{*}=0.5$. The performance measures are the average point estimate of the polychoric correlation coefficient, $\widehat{\rho}_{N}$, average bias ($\widehat{\rho}_{N}-\rho_{*}$), the standard deviation of the $\widehat{\rho}_{N}$ (“StDev”), the estimator’s coverage with respect to the true $\rho_{*}$ at significance level $\alpha=0.05$, as well as the length of the estimator’s confidence interval, again at level $\alpha=0.05$.
@@ 6.2 Polychoric correlation matrix
Let there be $q$ observed ordinal random variables where a latent variable underlies each ordinal variable. In accordance with the multivariate polychoric model (e.g., Muthén,, 1984), the latent variables are assumed to jointly follow a $q$-dimensional normal distribution with mean zero and a covariance matrix with unit diagonal elements so that the covariance matrix is a correlation matrix. Each individual latent variable is discretized to its corresponding observed ordinal variable akin to discretization process (1).
@@
For misspecification fraction $\varepsilon\in\{0,0.1,0.2\}$, we sample $N=1,000$ ordinal five-dimensional responses from this data generating process and use them to estimate the polychoric correlation matrix in Table 2 via ML and our robust estimator (again with tuning constant $c=1.6$). We repeat this procedure 1,000 times.
@@ 7.1 Background and study design
Arias et al., (2020) collect measurements of three Big Five traits in this way, namely extroversion, neuroticism, and conscientiousness. [^4] Arias et al., (2020) synonymously refer to neuroticism as emotional stability. Furthermore, in addition to the three listed traits, Arias et al., (2020) collect measurements of the trait dispositional optimism by using a different instrument, and another scale that is designed to not measure any construct. We do not consider these scales in this empirical demonstration. The sample that we shall use, Sample 1 in Arias et al., (2020), consists of $N=725$ online respondents who are all U.S. citizens, native English speakers, and tend to have relatively high levels of reported education (about 90% report to hold an undergraduate or higher degree). Concerned about respondent inattention in their data, Arias et al., (2020) construct a factor mixture model for detecting inattentive/careless participants. Their model crucially relies on response inconsistencies to polar opposite adjectives and is designed to primarily detect careless straightlining responding. They find that careless responding is a sizable problem in their data. Their model finds evidence of straightliners, and the authors conclude that if unaccounted for, they can substantially deteriorate the fit of theoretical models, produce spurious variance, and overall jeopardize the validity of research results.
Due to the suspected presence of careless respondents, we apply our proposed method to estimate the polychoric correlation coefficients between all $\binom{12}{2}=66$ unique item pairs in the neuroticism scale to obtain an estimate of the scale’s (polychoric) correlation matrix. The results of the remaining two scales are qualitatively similar and are reported in Appendix C. We estimate the polychoric correlation matrix twice: via classic maximum likelihood and via our proposed robust alternative, with tuning parameter $c=1.6$. The results remain qualitatively similar for different finite choices of $c$.
@@ 7.2 Results
Table 3: Parameter estimates with standard errors (SEs) for the correlation between the neuroticism adjective pair “envious” and “not envious” in the data of Arias et al., (2020), using Pearson sample correlation, ML, and robust estimator with tuning constant $c=1.6$. Each adjective item has five Likert-type answer categories. The sample correlation coefficient does not model thresholds, hence no estimates can be reported for them.
@@
1 $>1,000$ $>1,000$ 11.82 1.14 0.65
2 $>1,000$ 10.07 0.80 0.90 1.42
3 5.48 0.65 0.99 0.80 77.14
4 0.88 0.92 0.61 12.66 $>1,000$
5 0.89 0.88 36.01 $>1,000$ $>1,000$
(c) DPRs $\widehat{f}_{N}(x,y)\big{/}p_{xy}(\widehat{\bm{\theta}}_{N})$
@@
Table 4: Empirical relative frequency (top), estimated response probability (center), and decentralized Pearson residual (bottom) of each response $(x,y)$ for the “not envious” ($X$)–“envious” ($Y$) item pair in the measurements of Arias et al., (2020) of the neuroticism scale. Estimate $\widehat{\bm{\theta}}_{N}$ was computed with tuning constant $c=1.6$. The complete DPR values are provided in Table C.5 in the appendix.
@@
To study the potential presence of careless responses in each contingency table cell $(x,y)$ for item pair “envious” and “not envious”, Table 4 lists the DPRs at the robust estimate, alongside the associated model probabilities and empirical relative frequencies. [^5] A visualization of Table 4 is provided in Figure C.5 in the appendix. A total of six cells have extremely large DPR values of higher than 1,000, and, in addition, five cells have a DPR of higher than 10, and one cell has a DPR of higher than 5. Such DPR values are far away from ideal value 1 at which the model would fit perfectly, thereby indicating a poor fit of the polychoric model for such responses. It stands out that all such poorly fitted cells are those whose responses might be viewed as inconsistent. Indeed, response cells $(x,y)=(1,1),(1,2),(2,1),(2,2)$ indicate that a participant reports that neither “envious” nor “not envious” characterizes them accurately, which are mutually contradicting responses, while for response cells $(x,y)=(4,4),(4,5),(5,4),(5,5)$ both adjectives characterize them accurately, which is again contradicting. As discussed previously, such responses might due to careless responding. The robust estimator suggests that such responses cannot be fitted well by the polychoric model and subsequently downweighs their influence in the estimation procedure by mapping their Pearson residual with the linear part of the discrepancy function $\varphi(\cdot)$ in (11). Notably, also cells $(x,y)=(1,3),(3,1),(3,5),(5,3)$ are poorly fitted. These responses report (dis)agreement to one opposite adjective, while being neutral about the other opposite. It is beyond the scope of this paper to assess whether such response patterns are also indicative of careless responding, but the robust estimator suggests that such responses at least cannot be fitted well by the polychoric model with the data of Arias et al., (2020).
@@
Overall, leveraging our robust estimator, we find evidence for the presence of careless respondents in the data of Arias et al., (2020). While they substantially affect the correlation estimate of ML, amounting to about $-0.62$, which is much weaker than one would expect for polar opposite items, our robust estimator can withstand their influence with an estimate of about $-0.93$ and also help identify them through unreasonably large DPR values.
@@ 8.2 Simulation study
Figure 8: Boxplot visualization of the bias of the robust and ML estimator for the polychoric correlation coefficient, $\widehat{\rho}_{N}-\rho_{G}$, under distributional misspecification through a Clayton copula with correlation $\rho_{G}=0.9$ (left) and $\rho_{G}=0.3$ (right), across 1,000 repetitions. The tuning constant of the robust estimator is set to $c=1.6$.
@@ 9 Discussion and conclusion
We consider a situation where the polychoric correlation model is potentially misspecified for a subset of responses in a sample, that is, a set of uninformative observations not generated by a latent standard normal distribution. This model misspecification framework, called partial misspecification here, stems from the robust statistics literature, and a relevant special case is that of careless respondents in questionnaire studies. We demonstrate that maximum likelihood estimation is highly susceptible to the presence of such uninformative responses, resulting in possibly large estimation biases. As a remedy, we propose an estimator based on Welz, (2024) that is designed to be robust to partial model misspecification. Our estimator generalizes maximum likelihood estimation, does not make any assumption on the magnitude or type of potential misspecification, comes at no additional computational cost, and possesses attractive statistical guarantees such as asymptotic normality. It furthermore allows to pinpoint the sources of potential model misspecification through the notion of decentralized Pearson residuals (DPRs). Each possible response option is assigned a DPR, where values substantially larger than the ideal value 1 imply that the response in question cannot be fitted well by the polychoric correlation model. In addition, the methodology proposed in this paper is implemented in the free open source package robcat for the statistical programming environment R and is publicly available at https://github.com/mwelz/robcat.
@@
However, our estimator depends on a user-specified choice of a tuning constant $c$, which governs a tradeoff between robustness and efficiency. While simulation experiments suggest that the choice $c=1.6$ provides a good tradeoff and estimates do not change considerably for other finite choices of $c$, a detailed investigation on this tuning constant needs to be carried out in future work.
@@ A.1 Main theorem
For fixed tuning constant $c\geq 1$, let
$w(z)=\mathds{1}\left\{z\in[0,c]\right\}+c\mathds{1}\left\{z>c\right\}/z\quad\mathrm{for}\ z\geq 0,$
with first derivative
$w^{\prime}(z)=0\mathds{1}\left\{z\in[0,c]\right\}-c\mathds{1}\left\{z>c\right\}/z^{2}.$
Define the $d$-dimensional gradient of $\log\left(p_{xy}(\bm{\theta})\right)$ for response $(x,y)\in\mathcal{X}\times\mathcal{Y}$ at parameter $\bm{\theta}\in\bm{\Theta}$ as
$\bm{s}_{xy}(\bm{\theta})=\frac{1}{p_{xy}(\bm{\theta})}\left(\frac{\partial}{\partial\bm{\theta}}\ p_{xy}(\bm{\theta})\right),$
where one can write for the gradient of $p_{xy}(\bm{\theta})$
$\frac{\partial}{\partial\bm{\theta}}\ p_{xy}(\bm{\theta})=\frac{\partial}{\partial\bm{\theta}}\Phi_{2}\left(a_{x},b_{y};\rho\right)-\frac{\partial}{\partial\bm{\theta}}\Phi_{2}\left(a_{x-1},b_{y};\rho\right)-\frac{\partial}{\partial\bm{\theta}}\Phi_{2}\left(a_{x},b_{y-1};\rho\right)+\frac{\partial}{\partial\bm{\theta}}\Phi_{2}\left(a_{x-1},b_{y-1};\rho\right),$
see, e.g., Olsson, (1979, equation 4), and define the $d\times d$ Hessian matrix of $\log\left(p_{xy}(\bm{\theta})\right)$ as
$\bm{Q}_{xy}(\bm{\theta})=\frac{1}{p_{xy}(\bm{\theta})}\left(\frac{\partial^{2}}{\partial{\bm{\theta}}\partial{\bm{\theta}}^{\top}}p_{xy}(\bm{\theta})\right)-\bm{s}_{xy}(\bm{\theta})\bm{s}_{xy}(\bm{\theta})^{\top}.$ (A.1)
We derive closed-form expressions of all components in $\bm{Q}_{xy}(\bm{\theta})$ in Subsection A.3. In addition, for $K=K_{X}\cdot K_{Y}$, the total number of contingency table cells and $d$-dimensional vectors
$\bm{w}_{xy}\left(\bm{\theta}\right)=\bm{s}_{xy}(\bm{\theta})\mathds{1}\left\{\frac{f_{\varepsilon}\left(x,y\right)}{p_{xy}(\bm{\theta})}\in[0,c]\right\},$
@@
that row-binds all $K$ vectors $\bm{s}_{xy}(\bm{\theta})$ multiplied by an indicator that takes value 1 when associated population DPR is in the MLE-part of the function $\varphi(\cdot)$ in (11) and 0 otherwise. In similar fashion, define the $K$-dimensional vector
@@
Under certain regularity conditions that do not restrict the degree or type of possible misspecification of the polychoric model beyond $\varepsilon\in[0,0.5)$, when $N\to\infty$ it holds true that
@@
$\bm{M}\left(\bm{\theta}\right)$ $=\sum_{x\in\mathcal{X}}\sum_{y\in\mathcal{Y}}f_{\varepsilon}\left(x,y\right)\Bigg{(}{w^{\prime}}\Bigg{(}\frac{f_{\varepsilon}\left(x,y\right)}{p_{xy}(\bm{\theta})}\Bigg{)}\frac{f_{\varepsilon}\left(x,y\right)}{p_{xy}(\bm{\theta})}\bm{s}_{xy}(\bm{\theta})\bm{s}_{xy}(\bm{\theta})^{\top}-{w}\Bigg{(}\frac{f_{\varepsilon}\left(x,y\right)}{p_{xy}(\bm{\theta})}\Bigg{)}\bm{Q}_{xy}(\bm{\theta})\Bigg{)}$
@@
We list and discuss the regularity conditions in Subsection A.2.
@@
A.2 Assumptions
In the following, we list and discuss a set of assumptions that constitute the regularity conditions in Theorem 1. These assumptions are based on Welz, (2024).
Assumption Set A.
Suppose that the following assumptions hold true.
  1. A.1
    $c\in[1,+\infty]$,
  2. A.2
    $\bm{\Theta}\subset\mathbb{R}^{d}$ is compact, where $d=K_{X}+K_{Y}-1$ denotes the number of parameters in the polychoric model,
  3. A.3
    $\bm{\theta}_{0}=\arg\min_{\bm{\theta}\in\bm{\Theta}}L(\bm{\theta},f_{\varepsilon})$ is a unique minimum,
  4. A.4
    #$\#\big{\{}(x,y)\in\mathcal{X}\times\mathcal{Y}:f_{\varepsilon}\left(x,y\right)>0\big{\}}>d$,
  5. A.5
    $\frac{f_{\varepsilon}\left(x,y\right)}{p_{xy}(\bm{\theta}_{0})}\neq c$ for all $(x,y)\in\mathcal{X}\times\mathcal{Y}$.
Assumptions A.1A.3 are required for consistency of $\widehat{\bm{\theta}}_{N}$, whereas all assumptions are required for its asymptotic normality in Theorem 1. Under these assumptions, the theorem follows immediately by combining Theorems 1 and 2 in Welz, (2024).iiiThe general framework of Welz, (2024) entertains a total of nine assumptions instead of the five assumptions listed here. The four remaining assumptions are automatically satisfied in the polychoric model and are therefore omitted.
Assumption A.1 ensures that the discrepancy function $\varphi(\cdot)$ exhibits meaningful behavior when evaluated at DPRs, such as the ideal residual value 1 being included in the interval $[0,c]$. Compactness of the parameter space (Assumption A.2) is primitive and required for a technicality when proving consistency of $\widehat{\bm{\theta}}_{N}$ (see the proof of Theorem 1 in Welz,, 2024, for details). Uniqueness of a global minimum (Assumption A.3) is a common assumption in the literature on $M$-type estimators (e.g., Chapter 5.2 in Van der Vaart,, 1998). Assumption A.4 imposes that the number of positive response population probabilities is strictly larger than the number of model parameters. In other words, there must be more sources of variation (populated contingency table cells) than parameters. One may view this assumption as a rank condition that ensures invertibility of the Hessian matrix of the minimization problem (9), and is required to prevent rank deficiency of the asymptotic covariance matrix of estimator $\widehat{\bm{\theta}}_{N}$. Since $\widehat{f}_{N}(x,y)>0$ implies $f_{\varepsilon}\left(x,y\right)>0$, this assumption can be empirically verified. Finally, Assumption A.5 requires that no DPR at the global minimum is equal to tuning constant $c$. This is a primitive condition that is required for the loss function to be twice differentiable at $\bm{\theta}_{0}$, which is a requirement for the existence of the estimator’s asymptotic covariance matrix.
Finally, observe that no assumption is made that restricts the degree or type of potential misspecification of the polychoric model.
A.3 Expressions of first and second order derivatives
This section presents closed-form expressions of all components of matrix $\bm{Q}_{xy}(\bm{\theta})$ in (A.1).
A.3.1 First order terms

### v2

@@ 6.1 Individual polychoric correlation coefficient
For misspecification fraction $\varepsilon\in\{0,0.1,0.2\}$, we sample $N=1,000$ ordinal responses from this data generating process and estimate the true parameter $\bm{\theta}_{*}$ with the MLE (Olsson,, 1979), our proposed estimator with tuning constant set to $c=0.6$, and, for comparison, the Pearson sample correlation calculated on the integer-valued responses. We repeat this procedure for 1,000 simulated datasets. As performance measures, we calculate the average bias, standard deviation across repetitions, coverage, and length of confidence intervals at significance level $\alpha=0.05$. The coverage is defined as proportion of $(1-\alpha)$-th confidence intervals $[\widehat{\rho}_{N}\mp q_{1-\alpha/2}\cdot\text{SE}(\widehat{\rho}_{N})]$ that contain the true $\rho_{*}$, where $q_{1-\alpha/2}$ is the $(1-\alpha/2)$-th quantile of the standard normal distribution and $\text{SE}(\widehat{\rho}_{N})$ is the standard error of $\widehat{\rho}_{N}$, which is constructed using the limit theory developed in Theorem 1 in Appendix A. The length of a confidence interval is given by $2\cdot q_{1-\alpha/2}\cdot\text{SE}(\widehat{\rho}_{N})$.
Figure 4: Boxplot visualization of the bias of three estimators of the polychoric correlation coefficient, $\widehat{\rho}_{N}-\rho_{*}$, for various degrees of misspecification of the polychoric model across 1,000 replications. The estimators are the proposed robust estimator with $c=0.6$ (left), the MLE of Olsson, (1979) (center), and the Pearson sample correlation coefficient (right).
@@
Table 1: Performance measures of of the three estimators, Pearson sample correlation, the MLE, and the proposed robust estimator with $c=0.6$, for various misspecification fractions across 1,000 simulated datasets. The true polychoric correlation coefficient is $\rho_{*}=0.5$. The performance measures are the average point estimate of the polychoric correlation coefficient, $\widehat{\rho}_{N}$, average bias ($\widehat{\rho}_{N}-\rho_{*}$), the standard deviation of the $\widehat{\rho}_{N}$ (“StDev”), the estimator’s coverage with respect to the true $\rho_{*}$ at significance level $\alpha=0.05$, as well as the length of the estimator’s confidence interval, again at level $\alpha=0.05$.
@@ 6.2 Polychoric correlation matrix
Let there be $q$ observed ordinal random variables and assume that a latent variable underlies each ordinal variable. In accordance with the multivariate polychoric model (e.g., Muthén,, 1984), the latent variables are assumed to jointly follow a $q$-dimensional normal distribution with mean zero and a covariance matrix with unit diagonal elements so that the covariance matrix is a correlation matrix. Each individual latent variable is discretized to its corresponding observed ordinal variable akin to discretization process (1).
@@
For misspecification fraction $\varepsilon\in\{0,0.1,0.2\}$, we sample $N=1,000$ ordinal five-dimensional responses from this data generating process and use them to estimate the polychoric correlation matrix in Table 2 via ML and our robust estimator (again with tuning constant $c=0.6$). We repeat this procedure 1,000 times.
@@ 7.1 Background and study design
Arias et al., (2020) collect measurements of three Big Five traits in this way, namely extroversion, neuroticism, and conscientiousness. [^4] Arias et al., (2020) synonymously refer to neuroticism as emotional stability. Furthermore, in addition to the three listed traits, Arias et al., (2020) collect measurements of the trait dispositional optimism by using a different instrument, and another scale that is designed to not measure any construct. We do not consider these scales in this empirical demonstration. Furthermore, the data of Arias et al., (2020) has been made publicly available by the authors at https://osf.io/n6krb. The sample that we shall use, Sample 1 in Arias et al., (2020), consists of $N=725$ online respondents who are all U.S. citizens, native English speakers, and tend to have relatively high levels of reported education (about 90% report to hold an undergraduate or higher degree). Concerned about respondent inattention in their data, Arias et al., (2020) construct a factor mixture model for detecting inattentive/careless participants. Their model crucially relies on response inconsistencies to polar opposite adjectives and is designed to primarily detect careless straightlining responding. They find that careless responding is a sizable problem in their data. Their model finds evidence of straightliners, and the authors conclude that if unaccounted for, they can substantially deteriorate the fit of theoretical models, produce spurious variance, and overall jeopardize the validity of research results.
Due to the suspected presence of careless respondents, we apply our proposed method to estimate the polychoric correlation coefficients between all $\binom{12}{2}=66$ unique item pairs in the neuroticism scale to obtain an estimate of the scale’s (polychoric) correlation matrix. The results of the remaining two scales are qualitatively similar and are reported in Appendix C. We estimate the polychoric correlation matrix twice: via classic maximum likelihood and via our proposed robust alternative, with tuning parameter $c=0.6$. The results remain qualitatively similar for different finite choices of $c$.
@@ 7.2 Results
Table 3: Parameter estimates with standard errors (SEs) for the correlation between the neuroticism adjective pair “envious” and “not envious” in the data of Arias et al., (2020), using Pearson sample correlation, ML, and robust estimator with tuning constant $c=0.6$. Each adjective item has five Likert-type answer categories. The sample correlation coefficient does not model thresholds, hence no estimates can be reported for them.
@@
1 $>1,000$ $>1,000$ 10.82 0.14 $-0.35$
2 $>1,000$ 9.07 $-0.20$ $-0.10$ 0.42
3 4.48 $-0.35$ $-0.01$ $-0.20$ 76.14
4 $-0.12$ $-0.08$ $-0.39$ 10.66 $>1,000$
5 $-0.11$ $-0.12$ 35.01 $>1,000$ $>1,000$
(c) PRs $\widehat{f}_{N}(x,y)\big{/}p_{xy}(\widehat{\bm{\theta}}_{N})-1$
@@
Table 4: Empirical relative frequency (top), estimated response probability (center), and Pearson residual (bottom) of each response $(x,y)$ for the “not envious” ($X$)–“envious” ($Y$) item pair in the measurements of Arias et al., (2020) of the neuroticism scale. Estimate $\widehat{\bm{\theta}}_{N}$ was computed with tuning constant $c=0.6$. The complete PR values are provided in Table C.5 in the appendix.
@@
To study the potential presence of careless responses in each contingency table cell $(x,y)$ for item pair “envious” and “not envious”, Table 4 lists the PRs at the robust estimate, alongside the associated model probabilities and empirical relative frequencies. [^5] A visualization of Table 4 is provided in Figure C.5 in the appendix. A total of six cells have extremely large PR values of higher than 1,000, and, in addition, five cells have a PR of higher than 9, and one cell has a PR of higher than 4. Such PR values are far away from ideal value 0 at which the model would fit perfectly, thereby indicating a poor fit of the polychoric model for such responses. It stands out that all such poorly fitted cells are those whose responses might be viewed as inconsistent. Indeed, response cells $(x,y)=(1,1),(1,2),(2,1),(2,2)$ indicate that a participant reports that neither “envious” nor “not envious” characterizes them accurately, which are mutually contradicting responses, while for response cells $(x,y)=(4,4),(4,5),(5,4),(5,5)$ both adjectives characterize them accurately, which is again contradicting. As discussed previously, such responses might due to careless responding. The robust estimator suggests that such responses cannot be fitted well by the polychoric model and subsequently downweighs their influence in the estimation procedure by mapping their Pearson residual with the linear part of the discrepancy function $\varphi(\cdot)$ in (11). Notably, also cells $(x,y)=(1,3),(3,1),(3,5),(5,3)$ are poorly fitted. These responses report (dis)agreement to one opposite adjective, while being neutral about the other opposite. It is beyond the scope of this paper to assess whether such response patterns are also indicative of careless responding, but the robust estimator suggests that such responses at least cannot be fitted well by the polychoric model with the data of Arias et al., (2020).
@@
Overall, leveraging our robust estimator, we find evidence for the presence of careless respondents in the data of Arias et al., (2020). While they substantially affect the correlation estimate of ML, amounting to about $-0.62$, which is much weaker than one would expect for polar opposite items, our robust estimator can withstand their influence with an estimate of about $-0.93$ and also help identify them through unreasonably large PR values.
@@ 8.2 Simulation study
Figure 8: Boxplot visualization of the bias of the robust and ML estimator for the polychoric correlation coefficient, $\widehat{\rho}_{N}-\rho_{G}$, under distributional misspecification through a Clayton copula with correlation $\rho_{G}=0.9$ (left) and $\rho_{G}=0.3$ (right), across 1,000 repetitions. The tuning constant of the robust estimator is set to $c=0.6$.
@@ 9 Discussion and conclusion
We consider a situation where the polychoric correlation model is potentially misspecified for a subset of responses in a sample, that is, a set of uninformative observations not generated by a latent standard normal distribution. This model misspecification framework, called partial misspecification here, stems from the robust statistics literature, and a relevant special case is that of careless respondents in questionnaire studies. We demonstrate that maximum likelihood estimation is highly susceptible to the presence of such uninformative responses, resulting in possibly large estimation biases. As a remedy, we propose an estimator based on Welz, (2024) that is designed to be robust to partial model misspecification. Our estimator generalizes maximum likelihood estimation, does not make any assumption on the magnitude or type of potential misspecification, comes at no additional computational cost, and possesses attractive statistical guarantees such as asymptotic normality. It furthermore allows to pinpoint the sources of potential model misspecification through the notion of Pearson residuals (PRs). Each possible response option is assigned a PR, where values substantially larger than the ideal value 0 imply that the response in question cannot be fitted well by the polychoric correlation model. In addition, the methodology proposed in this paper is implemented in the free open source package robcat for the statistical programming environment R and is publicly available at https://github.com/mwelz/robcat.
@@
However, our estimator depends on a user-specified choice of a tuning constant $c$, which governs a tradeoff between robustness and efficiency in case of misspecification. While simulation experiments suggest that the choice $c=0.6$ provides a good tradeoff and estimates do not change considerably for other finite choices of $c$, a detailed investigation on this tuning constant needs to be carried out in future work. As an alternative, one could consider other discrepancy functions that do not depend on tuning constants, like the ones discussed in Welz, (2024).
@@ A.1 Main theorem
Define the $d$-dimensional gradient vector of $\log\left(p_{xy}(\bm{\theta})\right)$ for response $(x,y)\in\mathcal{X}\times\mathcal{Y}$ at parameter $\bm{\theta}\in\bm{\Theta}$ as
$\bm{s}_{xy}(\bm{\theta})=\frac{1}{p_{xy}(\bm{\theta})}\left(\frac{\partial}{\partial\bm{\theta}}\ p_{xy}(\bm{\theta})\right).$
Further, given tuning constant $c\in[0,\infty]$, define the $d\times d$ matrix
$\bm{A}_{xy}(\bm{\theta})=\Bigg{(}\frac{f_{\varepsilon}\left(x,y\right)}{p_{xy}(\bm{\theta})}\mathds{1}\left\{\frac{f_{\varepsilon}\left(x,y\right)}{p_{xy}(\bm{\theta})}-1\in[-1,c]\right\}+(c+1)\mathds{1}\left\{\frac{f_{\varepsilon}\left(x,y\right)}{p_{xy}(\bm{\theta})}-1\ >c\right\}\Bigg{)}\frac{\partial^{2}}{\partial{\bm{\theta}}\partial{\bm{\theta}}^{\top}}p_{xy}(\bm{\theta})$
for $(x,y)\in\mathcal{X}\times\mathcal{Y}$ and $\bm{\theta}\in\bm{\Theta}$. We derive closed-form expressions of the gradient and Hessian matrix of $p_{xy}(\bm{\theta})$ in Subsection A.2. In addition, for $K=K_{X}\cdot K_{Y}$ the total number of contingency table cells and $d$-dimensional vectors
$\bm{w}_{xy}\left(\bm{\theta}\right)=\bm{s}_{xy}(\bm{\theta})\mathds{1}\left\{\frac{f_{\varepsilon}\left(x,y\right)}{p_{xy}(\bm{\theta})}-1\in[-1,c]\right\},$
@@
that row-binds all $K$ vectors $\bm{s}_{xy}(\bm{\theta})$ multiplied by an indicator that takes value 1 when associated population Pearson residual is in the MLE-part of the function $\varphi(\cdot)$ in (11) and 0 otherwise. In similar fashion, define the $K$-dimensional vector
@@
For $c\in[0,\infty]$ the tuning constant in the discrepancy function, assume that $\frac{f_{\varepsilon}\left(x,y\right)}{p_{xy}(\bm{\theta}_{0})}-1\neq c$ for any $(x,y)\in\mathcal{X}\times\mathcal{Y}$. Then, under certain regularity conditions that do not restrict the degree or type of possible misspecification of the polychoric model beyond $\varepsilon\in[0,0.5)$, when $N\to\infty$ it holds true that
@@
$\bm{M}\left(\bm{\theta}\right)$ $=\sum_{x\in\mathcal{X}}\sum_{y\in\mathcal{Y}}\Bigg{[}\mathds{1}\left\{\frac{f_{\varepsilon}\left(x,y\right)}{p_{xy}(\bm{\theta})}-1\in[-1,c]\right\}\bm{s}_{xy}(\bm{\theta})\bm{s}_{xy}(\bm{\theta})^{\top}-\bm{A}_{xy}(\bm{\theta})\Bigg{]}$
@@
The regularity conditions are presented and discussed in Welz, (2024). We stress again that no assumption is made that would restrict the degree or type of potential misspecification of the polychoric model.
@@
For the choice $c=0$, note that the condition $\frac{f_{\varepsilon}\left(x,y\right)}{p_{xy}(\bm{\theta}_{0})}-1\neq c$ in Theorem 1 rules out the zero-misspecification case ($\varepsilon=0$). Indeed, Welz, (2024) shows that for $c=0$, the estimator is not asymptotically Gaussian if the model is correctly specified. We therefore recommend to choose $c>0$ (see Section 5.3 for a discussion).
A.2 Expressions of first and second order derivatives
This section presents closed-form expressions of all components of the gradient and Hessian matrix of the probability mass function of the polychoric model, $p_{xy}(\bm{\theta})$.
A.2.1 First order terms